1) How many grams of hydrocortisone should be used in preparing 105 suppositories, each weighing 2 g and containing 1% of hydrocortisone? (use one decimal place)

2) Express the following concentration as a ratio strength. Your answer will be 1:X. Determine the value for Xand enter ONLY the value of X in the answer box. 16 mg of active ingredient in 16 mL of solution

3) How many grams of sodium fluoride should be added to 145000 liters of drinking water containing 0.75 ppm of sodium fluoride to provide a recommended concentration of 1.75 ppm? (round to a whole number)
Answer #1

Solution:

1) 1% of hydrocortisone in 2 g can be expressed by mass percent formula,

Mass percent = (Mass of hydrocartisone / Total mass) x 100

1 = (Mass of hydrocartisone /2 g ) x 100

Mass of hydrocartisone = 2 g / 100 = 0.02 g

Thus, total mass of hydrocartisone in 105 suppositories

= 105 x 0.02 g = 2.1 g

2) Ratio strength is expresed as,

Ratio strength = mg /mL

= 16 mg / 16 mL = 1 / 1

Thus, X =1